How many yards of material from a 24 yard length of cloth remain after 3 pieces each of 3 1/2 yard long and 5 pieces of 2 1/4 yard long are remove
Answer by edjones(8007)

3 1/2=7/2
2 1/4=9/4
.
7/2 * 3 = 21/2 = 42/4
9/4 * 5 = 45/4
.
42/4 + 45/4 = 87/4 = 21 3/4
23 4/4 - 21 3/4 = 2 1/4 yards remain
.
Ed